# Steering AI-Driven Personalization of Scientific Text for General Audiences

### v1

@@ Abstract.
Digital media platforms (e.g., social media, science blogs) offer opportunities to communicate scientific content to general audiences at scale. However, these audiences vary in their scientific expertise, literacy levels, and personal backgrounds, making effective science communication challenging. To address this challenge, we designed TranSlider, an AI-powered tool that generates personalized translations of scientific text based on individual user profiles (e.g., hobbies, location, and education). Our tool features an interactive slider that allows users to steer the degree of personalization from 0 (weakly relatable) to 100 (strongly relatable), leveraging LLMs to generate the translations with given degrees. Through an exploratory study with 15 participants, we investigated both the utility of these AI-personalized translations and how interactive reading features influenced users’ understanding and reading experiences. We found that participants who preferred higher degrees of personalization appreciated the relatable and contextual translations, while those who preferred lower degrees valued concise translations with subtle contextualization. Furthermore, participants reported the compounding effect of multiple translations on their understanding of scientific content. Given these findings, we discuss several implications of AI-personalized translation tools in facilitating communication in collaborative contexts.
AI Personalization, Generative AI, Large language models, Science communication, Steering AI
ccs: Human-centered computing Interactive systems and toolsccs: Computing methodologies Natural language generation
@@ 1. Introduction
Science communication refers to the process of “communicating complex scientific information with general audiences to improve public awareness, interest, and understanding of science” (Burns et al., 2003). Both academic and industry professionals strive to make their ideas more accessible. As a result, science communication platforms, such as science blogs (Hogan, 2024) and science magazines (Times, 2024), have been translating scientific information into more accessible formats for decades. However, a single version of a text often falls short of meeting the diverse needs of a general audience (August et al., 2024). For instance, a journalist with an interest in quantum physics would have vastly different context and comprehension needs than a college student majoring in physics.
Scientific research articles often employ domain-specific language, including jargon and complex sentence structures, which can pose significant barriers to comprehension for general audiences. Large language models (LLMs) offer the capability to personalize content to various user contexts, transforming both the style (e.g.,  casual to formal (Das et al., 2023)) and content (e.g.,  simplifying complex ideas (August et al., 2024)). This adaptive personalization creates opportunities for enhancing audience engagement and understanding through more effective content dissemination (Kim et al., 2024). For example, LLMs can tailor the tone and style of complex concepts like scientific information to align with individual language preferences, thereby improving readability and comprehension across various fields (August et al., 2023; Das et al., 2023; Ding et al., 2023). Notably, AI can achieve this personalization at scale (i.e.,  AI-scalable personalization), generating multiple tailored variations of content for diverse audiences (August et al., 2024; Riedl, 2010; Kim et al., 2024). This scalability positions LLMs as powerful tools in advancing inclusive and effective science communication (Razack et al., 2021).
Among the many strategies in science communication, analogies have proven effective in translating technical content into more accessible forms (August et al., 2020). For instance, the structure of the solar system is often used as an analogy to explain the structure of an atom. While such analogies can broaden understanding of scientific information, they are typically generalized and may not resonate equally well with all audiences, as individual comprehension is shaped by societal, cultural, educational, and personal backgrounds. For example, the space analogy assumes familiarity with the solar system (Kim et al., 2024). LLMs have demonstrated the ability to generate personalized analogies that can help people understand complex concepts and ideas within their own context (Ding et al., 2023). Building on this capability, we leverage LLMs to generate multiple personalized analogies tailored to individual readers’ contexts and comprehension levels.
@@
To answer these questions, we conducted a user study with 15 non-expert participants who used the tool to understand two scientific texts by exploring multiple personalized translations. We conducted post-session semi-structured interviews to elicit feedback.
Participants generally found the analogies useful in understanding the content. Some favored general analogies (e.g., a participant liked a construction analogy describing the body’s cells as building sites—low personalization), while others preferred detailed, personalized analogies (e.g., a participant appreciated a baking analogy to explain harmful emissions from lithium-ion batteries—high personalization). Participants noted that reading multiple translations with varied analogies allowed them to piece together a fuller understanding of the scientific text, correcting misunderstandings along the way. Overall, users found the tool beneficial for learning about unfamiliar topics but were cautious about the reliability of AI-generated content. Based on these findings, we discuss the need for HCI techniques to steer models towards human pluralistic preferences, rather than solely relying on machine learning methods that assume uniformity in human experiences.
In summary, our study has the following three contributions:
@@
    A novel slider-based interaction to enhance understanding of scientific content through the exploration of analogy-driven personalized translations
@@
    An investigation to examine the utility, benefits, and limitations of personalized translations in science communication
@@
    Implications for designing such AI-personalized translation tools and broader applications in collaborative and education contexts such as facilitating cross-disciplinary communication
@@ 2. Related Work
Below, we review prior work covering science communication on digital media platforms and personalization of language content for better science communication.
@@ 2.1. Science Communication on Digital Media Platforms
Science communication takes many forms in modern media for communicating complex scientific information to general audiences for improving public awareness, interest, and understanding of science (Burns et al., 2003). Traditional channels include museums (Bell, 2008), exhibitions (Long et al., 2021), and TV series  (Sagan et al., 1980), while web-based digital platforms have emerged as a powerful new stream for the broader public. For example, online videos have become an influential source for learning science, especially during the COVID-19 pandemic, when many people turned to digital resources for scientific information (Breslyn and Green, 2022). Social media platforms such as Reddit (e.g., r/science), Twitter (now X), Bluesky, and Mastodon have evolved into vibrant spaces where some scientists actively share their research, promote their publications, and engage in public discussions (Gero et al., 2021; Jones et al., 2019). Additionally, science blogs have been valuable channels for science communication, offering tailored narratives of complex scientific content for general audiences (Jarreau and Porter, 2018). One of the core benefits of science communication on these digital media is their scalability in reaching broader audiences (e.g.,  (Nguyen et al., 2024; Jones et al., 2019)). Recent studies show that more than half of college students actively use blogs as a learning resource (Head et al., 2017).
@@
Prior HCI/CSCW literature has uncovered several key challenges that scientists face in science communication: engagement, translation, and dissemination (Williams et al., 2022). First, scientists rarely participate in digital media platforms (Zhu and Purdam, 2017). This limited engagement may be associated with multiple factors, including conventional academic culture that has historically put scientists as neutral observers rather than active communicators (Martinez-Conde, 2016; Peters, 2013), and practical constraints such as time limitations and lack of communication training (Besley et al., 2015). While recent research shows a gradual shift towards more active public engagement by scientists (Brüggemann et al., 2020), the extent and impact of this change remain to be fully understood. Secondly, even for those scientists who engage with digital communication, translating complex scientific concepts into accessible content for general audiences is challenging (Rice and Giles, 2017). These challenges may be compounded by the varied characteristics of general audiences in digital media platforms, where readers have widely differing backgrounds, levels of scientific literacy, and preferences (Schäfer, 2017; Su et al., 2014; Rice and Giles, 2017). A single version of scientific content translation would not sufficiently serve this wide spectrum of general audiences, yet creating multiple versions tailored to different expertise levels would also be infeasible for individual scientists (Bucchi and Trench, 2014; Fischhoff, 2013). Lastly, such translation challenges can lead to broader concerns about dissemination, particularly regarding the potential for miscommunication and misinformation on social media platforms (Cook et al., 2018). Scientists may be hesitant to engage in digital communication because the dissemination of inaccurate interpretations or misrepresentations of their work could have unwanted consequences for public understanding and their professional reputation (Wu et al., 2019; Loroño-Leturiondo and Davies, 2018).
@@
Given these challenges, our study focuses on addressing the translation barrier in personalized science communication on digital media platforms. Specifically, we explore how to facilitate creating multiple translations of scientific content to effectively accommodate large-scale general audiences with varying levels of scientific background.
2.2. Personalization of Language Content for Better Science Communication
Prior research has explored various approaches to make complex science texts more accessible and comprehensible to broader audiences, from using plain language summaries (August et al., 2023) and language constructs like analogies (Kim et al., 2022) to using interactive techniques and systems (Chang et al., 2023; Fok et al., 2023). Across these works, limited explorations have been made in personalization of scientific translations to facilitate better understanding, which this paper explores.
Personalization has been explored across various forms of language content, including news articles (Adar et al., 2017), social media posts (Zhang et al., 2022), and novels (Kim et al., 2024), to better accommodate individual users’ needs and preferences. Prior research has demonstrated the effectiveness of such personalization across several domains (e.g., healthcare, journalism) through approaches such as using patients’ demographic and medical information to tailor content  (Dimarco et al., 2006), using users’ location data to make news content relevant (Adar et al., 2017), and using analogies to make measurements (e.g., distances, areas) comprehensible to a user’s context (Kim et al., 2016). For instance, employing familiar objects as an analogy (e.g., 28 lbs is about the weight of a microwave) to personalize the unit expressions (Hullman et al., 2018). In science communication contexts, past work has shown AI’s potential in large scale translation and dissemination through generating personalized summaries of scientific texts for general audiences (August et al., 2024). In this paper, we build on these ideas to explore using LLMs for generating and customizing scientific text explanations using user provided data for analogies.
While relying on users’ contextual information (e.g., location) delivers personalized outcomes, these systems typically operate autonomously making personalization decisions without the users’ direct input or control. This raises several potential concerns. First, there are privacy issues, as these systems require users to share personal data, often assuming their willingness to do so or even collecting such information without notice (implicitly) (e.g.,  location information based on IP address (Ma et al., 2024), personal profile based on search engine account (Hannak et al., 2013; Majumder and Shrivastava, 2013)). Second, users generally lack control over which specific data they share for personalization purposes. Additionally, individual users may have different preferences regarding how much of their profile information should be integrated with the tailored content. To address these issues in user agency and privacy control, we designed a steerable reading interface that enables users to control the degree of personalization with adjustable sliders as well to edit profile information directly for personalizing the translations to their context.
@@ 3. TranSlider: Design and Implementation
We designed and implemented TranSlider (Translate through Slider), an AI-powered tool to help individuals with no scientific backgrounds better understand scientific texts by personalizing them to their context using a slider interaction and user profiles.
Figure 1. The interface design of TranSlider. (A) The AI-personalized title of the science blog post appears here. (B) AI-personalized translations of science articles will appear here. (C) Users can navigate the slider to adjust their preferred degree for AI personalization of science blog intro. (D) For each personalized translation, a history item is generated and stacked in the history box. Users can revisit any previous history items to see the translations. (E) Users can click the edit profile button to see the Profile Edit window [II] to adjust their profile information. (F) The Finish button triggers a pop-up [III] to collect the user’s interest in reading the original paper. We use the response as a proxy measure for their interest in the content after reading the personalized translations.
@@ 3.1. Interface Design
The interface comprised three main components: the translation panel, the history box, and the edit profile dialog box (Figure 1). The header provided instructions on using the interface; these instructions were brief as participants received a detailed tutorial at the start of the study.

### v2

@@ Abstract.
Digital media platforms (e.g., science blogs) offer opportunities to communicate scientific content to general audiences at scale. However, these audiences vary in their scientific expertise, literacy levels, and personal backgrounds, making effective science communication challenging. To address this challenge, we designed TranSlider, an AI-powered tool that generates personalized translations of scientific text based on individual user profiles (e.g., hobbies, location, and education). Our tool features an interactive slider that allows users to steer the degree of personalization from 0 (weakly relatable) to 100 (strongly relatable), leveraging LLMs to generate the translations with chosen degrees. Through an exploratory study with 15 participants, we investigated both the utility of these AI-personalized translations and how interactive reading features influenced users’ understanding and reading experiences. We found that participants who preferred higher degrees of personalization appreciated the relatable and contextual translations, while those who preferred lower degrees valued concise translations with subtle contextualization. Furthermore, participants reported the compounding effect of multiple translations on their understanding of scientific content. Drawing on these findings, we discuss several implications for facilitating science communication and designing steerable interfaces to support human-AI alignment.
Human-AI alignment, Scalable personalization, Science communication, Steering AI, Translation
copyright: ccjournal: PACMHCIjournalyear: 2025journalvolume: 9journalnumber: 7article: CSCW479publicationmonth: 11doi: 10.1145/3757660ccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Human computer interaction (HCI)$c$s: Human-centered computing Empirical studies in HCIccs: Computing methodologies Natural language generation
@@ 1. Introduction
Science communication refers to the process of “communicating complex scientific information with general audiences [^1] We refer to “general audiences” as individuals who are non-scientists, coming from diverse educational backgrounds and domains such as arts, business, education, healthcare, and product design. to improve public awareness, interest, and understanding of science” (Burns et al., 2003). Both academic and industry professionals strive to make their ideas more accessible. As a result, science communication platforms, such as science blogs (Hogan, 2024) and science magazines (Times, 2024), have been translating scientific information into more accessible formats for decades. However, a single version of a text often falls short of meeting the diverse needs of a general audience (August et al., 2024). For instance, a journalist with an interest in quantum physics would have vastly different context and comprehension needs than a college student majoring in physics.
Scientific research articles often employ domain-specific language, including jargon and complex sentence structures, which can pose significant barriers to comprehension for general audiences. Language models offer the capability to personalize content to various user contexts, transforming both the style (e.g.,  casual to formal (Das et al., 2023) or romantic (Kim et al., 2019)) and content (e.g.,  simplifying complex ideas (August et al., 2024)). This adaptive personalization creates opportunities for enhancing audience engagement and understanding through more effective content dissemination (Kim et al., 2024). Notably, large language models (LLMs) can achieve this personalization at scale (i.e.,  AI-scalable personalization), generating multiple tailored variations of content for diverse audiences (August et al., 2024; Riedl, 2010; Kim et al., 2024). This scalability positions LLMs as powerful tools in advancing inclusive and effective science communication (Razack et al., 2021).
Among the many strategies in science communication from prior HCI/CSCW literature, analogies have proven effective in translating technical content into more accessible forms (August et al., 2020). For instance, the structure of the solar system is often used as an analogy to explain the structure of an atom. While such analogies can broaden understanding of scientific information, they are limited in number, can be generic, and may not resonate equally well with all audiences as individual comprehension is shaped by societal, cultural, educational, and personal backgrounds. For example, the space analogy assumes familiarity with the solar system (Kim et al., 2024). LLMs have demonstrated the ability to generate personalized analogies that can help people understand complex concepts and ideas within their own context (Ding et al., 2023). Building on this capability, we leverage LLMs to generate multiple personalized analogies tailored to individual readers’ contexts.
@@
To answer these questions, we conducted a user study with 15 non-expert participants who used the tool to understand two scientific texts by exploring multiple personalized translations. We conducted post-session semi-structured interviews to elicit feedback on their experience. We used thematic analysis on their qualitative responses and conducted descriptive quantitative analysis on tool usage logs to understand their exploration behavior (e.g., number of generated translations. range of explored personalization degrees).
Participants found the analogies useful in understanding the content. Some favored general analogies (e.g., a participant liked a construction analogy describing the body’s cells as building sites—low personalization), while others preferred detailed, personalized analogies (e.g., a participant appreciated a baking analogy to explain harmful emissions from lithium-ion batteries—high personalization). Participants noted that reading multiple translations with varied analogies allowed them to piece together a fuller understanding of the scientific text, correcting misunderstandings along the way. Participants found the tool beneficial for learning about unfamiliar topics but were cautious about the reliability of AI-generated translations. Based on these findings, we discuss implications for science communication and the need for interactive techniques to steer models towards human pluralistic preferences, rather than solely relying on machine learning approaches.
Overall, our contributions to the CSCW community include,
@@
    A novel slider-based LLM interaction to enhance understanding of scientific content through the exploration of analogy-driven personalized translations
@@
    An investigation to examine the utility, benefits, and limitations of AI-driven personalized translations in science communication
@@
    Implications for science communication and more broadly for cross-disciplinary communication, and designing interfaces for human-AI alignment.
@@ 2. Related Work
We review prior work covering science communication on digital media platforms, personalization in science communication, and existing interactive reading interfaces for scientific articles.
@@ 2.1. Science Communication on Digital Media Platforms
Science communication takes many forms in modern media for communicating complex scientific information to general audiences for improving public awareness, interest, and understanding of science (Burns et al., 2003). Traditional channels include museums (Bell, 2008), exhibitions (Long et al., 2021), and TV series (Sagan et al., 1980), while web-based digital platforms have emerged as a powerful new stream. For example, online videos have become an influential source for science learning, especially during the COVID-19 pandemic, when many people turned to digital resources for scientific information (Breslyn and Green, 2022). Social media platforms such as Reddit (e.g., r/science), Twitter (now X), Bluesky, and Mastodon offer interactive spaces for scientists to share research and engage with the public (Gero et al., 2021; Martin and MacDonald, 2020; Jones et al., 2019). Finally, science blogs have been valuable channels for explaining complex scientific content to general audiences (Jarreau and Porter, 2018). For example, more than half of college students actively use blogs as a learning resource (Head et al., 2017). A key advantage of science communication via digital media, a.k.a. online science communication (Williams et al., 2022), is its ability to reach a wider audience (e.g.,  (Nguyen et al., 2024; Jones et al., 2019)).
@@
Prior HCI/CSCW literature has uncovered several key challenges scientists face in online science communication: engagement, translation, and dissemination (Williams et al., 2022). First, scientists rarely participate in digital media platforms (Zhu and Purdam, 2017). This limited engagement is due to the conventional academic nature that has put scientists as neutral observers rather than active communicators (Martinez-Conde, 2016; Peters, 2013), and practical constraints such as limited time and lack of communication training (Besley et al., 2015). While recent research shows a gradual shift towards more active public engagement by scientists (Brüggemann et al., 2020), the extent and impact of this change remain to be fully understood. Secondly, even for scientists who engage with digital communication, translating complex scientific concepts into accessible language is challenging (Rice and Giles, 2017). These challenges may be compounded by the variety of readers with different backgrounds, levels of scientific literacy, and preferences (Schäfer, 2017; Su et al., 2014; Rice and Giles, 2017). Creating multiple versions tailored to different expertise levels is infeasible for individual scientists (Bucchi and Trench, 2014; Fischhoff, 2013). Lastly, translation challenges can entail broader concerns such as misinformation on social media platforms (Cook et al., 2018). The dissemination of inaccurate interpretations or misrepresentations of their work could have unwanted consequences for public understanding and their professional reputation (Wu et al., 2019; Loroño-Leturiondo and Davies, 2018).
@@
Given these challenges, this paper focuses on addressing the translation barrier in personalized science communication on digital media platforms. Specifically, we explore how to facilitate creating multiple translations of scientific content to effectively accommodate general audiences with varying levels of scientific background at scale.
2.2. Personalization in Science Communication: Tools and Approaches
In science communication, personalization is an approach that translates scientific content into contexts familiar to each individual, making it more interesting, accessible, and engaging for non-scientists (Anand and Mobasher, 2007). For example, when explaining biological concepts to individuals without a background in biology, saying “try not to drink soda and eat candy” may be more effective than “reduce glucose consumption”, as the former uses more familiar contexts while conveying the same information to the intended audience (Li et al., 2024). Given its potential effectiveness for science communication, researchers have explored various strategies for personalizing information across contexts.
[…]
These approaches predate the emergence of LLMs that could address these limitations: reliance on structured templates and databases. LLMs can transform text at scale without any databases, making it feasible to generate multiple translations on behalf of scientists. Recent HCI studies have explored the feasibility of LLMs in simplifying complex ideas (summarization) (August et al., 2023; Gu et al., 2024; Kim et al., 2015), retargeting concepts into another domain (contextualization) (Ding et al., 2023; Bao et al., 2025; Liu et al., 2025), and explaining content through user-familiar contexts (personalization) (August et al., 2024; Kim et al., 2024). Albeit in other domains, these studies show that LLMs can reduce the reliance on databases to compute contextual relevance and generate natural-sounding language expressions tailored to various user contexts. Building on these works, our study explores integrating contextual personalization with LLMs’ capability to generate analogies (Ding et al., 2023). We aim to design an LLM-powered tool that leverages readers’ contextual information to generate personalized analogies for scientific texts.
2.3. AI-Powered Reading Interfaces for Scientific Articles
Recent research has proposed AI-powered interactive reading interfaces for science articles. These systems summarize sections of the paper and help navigate within the article to efficiently find relevant information (August et al., 2023; Fok et al., 2023). For example, ScholarPhi assists users in quickly looking up definitions of nonce words and symbols defined elsewhere in the paper (Head et al., 2021). Researchers have also explored more free-form question-answering systems to help users identify relevant sections of a paper based on their queries (Zhao and Lee, 2020). However, these primarily cater to users already motivated to engage with scientific literature, such as healthcare stakeholders (e.g., patients (Kambhamettu et al., 2024)), students, and scientists.
While general-purpose tools like ChatGPT or AI-enabled PDF readers (e.g.,  Adobe Acrobat with built-in chatbots) can assist with reading scientific texts, they are not designed for scientific content, and require users to know what to ask. In contrast, our tool focuses on everyday readers with a casual interest in science—individuals who may lack the time, background, or motivation to actively engage with science articles. We position our tool as an AI-powered intermediary (Kim et al., 2024) that delivers accessible, engaging science communication through analogy-driven, personalized translations. Rather than expecting readers to extract meaning themselves, our tool meets them where they are by adapting content to their context.
3. Interaction Paradigms
3.1. Existing Design: Model-Initiative
Existing personalization systems predominantly employ a model-initiative design, where AI models automatically collect and process user data to deliver personalized content without explicit user intervention (Corrigan et al., 2014). While efficient, this interaction design presents two primary human-AI alignment issues, namely personality alignment and user privacy.
3.1.1. Personality Alignment Issues
Model-initiative approach faces the personality alignment issue (Zhu et al., 2025)i.e., the personalized outcomes are not always well aligned with users’ actual interests or needs (Zollo et al., 2025; Dudy et al., 2021). For example, consider an AI system that automatically personalizes results based on a user’s current location (e.g., Tokyo, Japan) when the user is actually interested in contextualized results of a different location (e.g., Bergen, Norway). In this case, the AI personalization is misaligned with the user’s actual interests. This alignment problem is compounded by the contextual nature of personalization preferences. Individuals may desire different levels of personalization depending on the context and motivation of their reading (August et al., 2024; Grunig, 1979; Leerink et al., 2024; Dudy et al., 2021). For example, someone might prefer highly personalized content when reading medical articles relevant to their health conditions (Lim et al., 2011), but relatively less or lower personalization when casually exploring scientific topics of general interest. To accommodate varying contextual needs for each individual at scale, the model-initiative approach would naturally require invasive data collection and inference about users’ contexts.
3.1.2. Privacy Issues
Model-initiative design assumes that AI models have access to users’ private data (Corrigan et al., 2014). However, from a user’s perspective, it is unclear how much information is gathered, what types of information is gathered, and when (Dudy et al., 2021). Furthermore, users typically have minimal control over their personal data (Lenhart et al., 2023). These privacy concerns become more problematic in the context of science communication, which aims to reach diverse, mass audiences (Burns et al., 2003). The model-initiative paradigm would require collecting and processing private data from the entire population of potential readers to provide personalized translations of scientific texts. This approach not only raises ethical concerns but also creates significant barriers to adoption, as many people would be reluctant to share their personal information.
3.2. Our Design: Controlled User-Initiative
We explore an alternative approach: user-initiative design. Existing LLM interfaces, such as ChatGPT (OpenAI, 2022) and Claude (Anthropic, 2024), provide prompt interfaces where users can specify all personalization parameters and their intentions. While a prompt-based approach maximizes user control, it presents challenges specifically for science communication: allowing unlimited user modification risks compromising the scientific accuracy crucial for effective information dissemination. High degrees of freedom could lead to the distortion of key ideas or inadvertently promote misinformation, even when users have no intention to mislead (Zhang and Horvát, 2025; Kim et al., 2024).
We propose a controlled user-initiative design where readers can explicitly adjust specific aspects for personalization. To minimize the risk of misinformation or hallucination, rather than allowing free-form prompt interactions that could introduce unexpected results, we constrained user inputs to two structured elements: i) steerable degrees of personalization using sliders, addressing the alignment issue, and ii) editable structured type of contextual information via user profiles with predefined fields to alleviate the privacy issue. This design philosophy aims to preserve privacy by minimizing detailed personal data collection and granting users control over input, while enabling contextually relevant personalization that aligns with their needs and preferences.
Figure 1. The interface design of TranSlider. A The AI-personalized title of the science blog post appears here. B AI-personalized translations of science articles will appear here. C Users can navigate the slider to adjust their preferred degree of AI personalization for a science article intro. D For each personalized translation, a history item is stacked in the history box. Users can revisit any previous history items to see the translations. E Users can click the edit profile button to see the Profile Edit window [II] to adjust their profile information. F The Finish button triggers a pop-up [III] to collect the user’s interest in reading the original paper. We use the response as a proxy measure for their interest in the content after reading the personalized translations.
@@ 4. TranSlider: Design and Implementation
With this design philosophy in mind, we designed and implemented TranSlider (Translate through Slider), an LLM-powered reading interface to help general audiences better understand scientific texts by generating personalized translations through a slider interaction and user profiles.
@@ 4.1. Interface Design
We designed our tool with three main components: the translation panel, the history box, and the edit profile button (Figure 1). The header provides instructions on using the interface; these instructions were brief as participants received a detailed tutorial at the start of the study.